OpenStudy (anonymous):

rationalize the denominator and simplify sq root a + 2 sq root y over sq root a - 2 sq root y

OpenStudy (yttrium):

\[\frac{ \sqrt{a} + 2\sqrt{y}}{ \sqrt{a} - 2\sqrt{y} } \frac{ \sqrt{a}+2\sqrt{y} }{ \sqrt{a}+2\sqrt{y} }\]

OpenStudy (yttrium):

That will be equal to: \[\frac{ (\sqrt{a}+2\sqrt{y})^2 }{ a-4y }\]

OpenStudy (yttrium):

which is same as\[\frac{ a+4\sqrt{ay}+4y }{ a-4y }\]